# Generative Models in Decision Making: A Survey

### v2

@@ 2.4 Difference with previous approaches
Reinforcement Learning (RL) and generative models are two distinct branches of machine learning that serve different purposes and tackle different challenges. The key difference lies in their primary objectives and the nature of the learning process [114], as show in Fig. 6.
The primary objective of RL is for an agent to learn an optimal policy that maximizes cumulative rewards through trial-and-error with the environment. The agent learns by taking actions, observing the resulting states, then adjusting current strategy according to the returns. On the other hand, generative models aims to understand and represent the latent data distribution in a given dataset [115]. These models are used to generate new samples resembling the training data, capturing inherent patterns and structures in the data.
The second one is the learning process. In RL, the agent learns through trial and error by exploring the environment, taking actions, and receiving rewards. As it gains experience, the agent refines its policy to make more effective decisions, ultimately maximizing its rewards. Meanwhile, generative models learn from a given dataset and analyze the data, attempting to estimate the probability distribution that generates the data, such as VAEs and GANs.
Last but not least, they differ in how they interact with environment [116]. The RL agent interacts with the environment iteratively, learning from feedback and adjusting its behavior without generating any data. In contrast, the generative model, once trained, is capable of producing entirely new data points resembling the original ones [117], thus actively generating rather than merely responding.
@@
In summary, reinforcement learning focuses on learning optimal actions to achieve specific goals [118], whereas generative models aim to model data distributions and generate new, similar samples. Both are crucial in machine learning but address different challenges.
@@
Du & Mordatch [119] EBM Controller EBM Imitation Learning Robot Control
@@
PixelCNN [85] AM Modeler Decoder Only Generation Structural Generation
PixelRNN [183] AM Modeler Decoder Only Generation Structural Generation
@@ 3 Taxonomy
Application. The last dimension highlights several representative applications where generative models have shown particular promise, including robot control, autonomous driving, gaming, structural generation, and optimization.
@@ 4.1 Generative Models as Controller
A controller aims to find a policy (a set of instructions or rules) that determines how to modify the system’s inputs to attain the desired behavior [186, 187]. The policy is typically generated through a process called control system design, which involves analyzing the system’s behavior, defining the desired outcomes, and then designing a controller to achieve those goals. The method used to generate the policy according to the system’s specific requirements and the design process’s objective. Generative models are able to model the decision-making process and produce diverse decision paths, making them well suited for generating policies in control systems.
Generative models are adept at capturing complex data distributions, making them particularly effective for learning from expert trajectories in decision-making tasks. When applied as controllers, these models can leverage offline data containing trajectories associated with high-reward values to extract and replicate optimal policies. In general, generative models facilitate the generation of control policies through two primary approaches, which we detail below.
(1) Generative models learn the policy from the state-action or observation-action pairs, as shown in (19):
$\theta=\arg\min_{\theta}\mathbb{E}_{(s_{i},a_{i})\sim D}L\left(\pi^{*}(s_{i})-\pi_{\theta}(s_{i})\right),$ (19)
where $L$ is a metric function.
(2) Generative models learn the policy from the decision trajectories $D={(s_{0},a_{0},s_{1},a_{1},\cdots)}$ as shown in (20):
$\theta=\arg\min_{\theta}\mathbb{E}_{(s_{0},a_{0},s_{1},a_{1},...)\sim D}L\left(\pi^{*}(s)-\pi_{\theta}(s)\right).$ (20)
The detailed explanations are as follows.
@@ 4.1.1 EBMs as Controller
EBMs serve as controllers by solving the Inverse Reinforcement Learning (IRL) problems, where the goal is to extract a control policy based on expert demonstrations. In this context, the energy-based model is used to learn the underlying reward structure of the expert’s behavior. Du & Mordatch [119] take a pioneering approach by using Markov Chain Monte Carlo (MCMC)-based EBM training to generate robotic hand trajectories, applying it to continuous neural networks. Building on this, EBIL [120] further refines the process by introducing a more straightforward solution to the IRL problems. Specifically, EBIL estimates the expert’s energy through score matching, which serves as a surrogate reward function. This reward function is then utilized to learn the policy through RL algorithms.
@@ 4.1.2 GANs as Controller
Generative Adversarial Imitation Learning (GAIL) [64] combines imitation learning with GANs to directly extract the policy from data. It seeks to find a saddle point $(\pi,D)$, representing a pair consisting of a policy $\pi$ and a discriminator $D$, by optimizing the following expression:
@@
where $(s,a)$ denotes the state-action pair, $\pi$ is the policy, $\pi_{E}$ represents the expert policy and $H$ is a regularizer for the policy. To overcome some of GAIL’s limitations, several extensions have been proposed. For example, InfoGAIL [125], which builds upon GAIL, aims to calculate the potential structure of expert demonstrations in an unsupervised manner. This method can both imitate complex behaviors and extract meaningful, interpretable representations from intricate behavioral data. Similarly, MGAIL [126] incorporates a forward model into GAIL, making the calculation differentiable and enabling policy training with fewer samples and interactions with the environment. In contrast, Fail-Safe Adversarial Generative Imitation Learning (FAGIL) [127], an extension of GAIL, introduces a safety layer to ensure the generated policy is safe, while also facilitating closed-form computation of probability densities and gradients. This method integrates end-to-end generative adversarial training with worst-case safety guarantees, making it well-suited for scenarios requiring both safety and performance. On the other hand, WGAIL [128] and IC-GAIL [129] address the issue of imperfect demonstrations in GAIL. WGAIL focuses on learning the weights of imperfect demonstrations without requiring substantial prior information, while IC-GAIL uses confidence levels to assess the quality of these demonstrations, improving learning efficiency in the presence of noisy or suboptimal expert data.
Moreover, there are other related approaches such as Adversarial Inverse Reinforcement Learning (AIRL) [130], which formulates policy learning as an adversarial reward learning problem, allowing for acquisition of policies in diverse environments. AugAIRL [131] enhances AIRL by adopting semantic rewards in the learning framework (see Fig. 7 for an overview), while WAIL [132] bridges inverse reinforcement learning with optimal transport, using regularized optimal transport for large-scale applications. Lastly, Model-based AIRL (MAIRL) [133] improves policy optimization by leveraging a self-attention dynamics model to make the computation graph end-to-end differentiable, and SC-AIRL [188] further reduces the need for extensive exploration by decomposing long-horizon tasks into subtasks, using shared rewards and critics across different subtasks.
@@ 4.1.3 VAEs as Controller
Variational Autoencoders (VAEs) serve as controllers by learning compact, structured latent spaces that encode high-dimensional environments and decision-making processes, which are directly used to generate control actions.
For example, Play-LMP [141] introduces a self-supervised approach to learn control from play behaviors, enabling the system to autonomously generate control policies. Building on this idea, TACO-RL [142] extends Play-LMP by adopting a hierarchical model, which allows for learning long-horizon policies from high-dimensional camera observations in agnostic tasks. OPAL [143] utilizes VAE for offline learning of primitive actions in imitation learning tasks. Similarly, Generalization Through Imitation (GTI) [139] develops a two-stage algorithm that leverages VAE for policy learning through imitation, enhancing generalization in complex environments. Finally, HULC [140] attempts to improve performance by integrating hierarchical decomposition in robot-controlled learning, utilizing a multimodal transformer encoder, leveraging discrete latent plans, and applying a self-supervised contrastive loss.
Additionally, MaskDP [144] introduces a masked autoencoding technique that enhances decision-making models by learning latent spaces in a scalable and generalizable manner, thereby improving performance in high-dimensional tasks (see Fig. 8). Similar to VAEs, this approach focuses on generating control policies by learning efficient latent representations, while additionally incorporating masking mechanisms to improve the model’s ability to generalize across tasks.
Figure 8: Illustration of MaskDP. In the pretraining phase, the model is trained to predict masked tokens. After pretraining, it can be applied to a variety of downstream tasks by using different masking patterns [144].
@@ 4.1.4 Normalizing Flows as Controller
Normalizing flows have garnered interest recently due to their capacity to learn complex probability distributions and perform efficient density estimation. While they are primarily used in generative modeling tasks, they can also be applied to decision-making problems control [75, 91]. In this context, normalizing flows can be leveraged to model and approximate the underlying distribution of the decision space[53, 89].
The natural use of Normalizing flows for controllers is behavioral cloning, where the behavior to reproduce is represented by the target distribution $\mu$. However, since the training only requires $\mu$ to be samplable and $\nu$ to have a defined density, one can train a ”reversed Normalization flows”, where $\nu$ represents a complex distribution defined by its density and $\mu$ is a simpler distribution, such as a normal distribution.
Normalizing Flow Policy (NF Policy) [148] integrates Normalizing Flow policies into the Soft Actor-Critic (SAC) framework, enabling the learning of more expressive policies. Fig. 9 illustrates the process of sampling actions from NF policies. Conservative Normalizing Flows (CNF) [149] makes use of Normalizing Flows as a conservative action encoder to learn policies in latent action spaces. First, the action encoder undergoes supervised pre-training on an offline dataset, and then a reinforcement learning-based policy model (serving as a controller in the latent space) is trained. Guided Flows [150] generates plan in the offline reinforcement learning setting by training Continuous Normalizing Flows based on regressing vector fields via integrating classifier free guidance.
Figure 9: A sampled action from NF policies [148].
@@ 4.1.5 Diffusion Models as Controller
Diffusion models, characterized by their iterative refinement and controlled noise perturbation processes, have proven to be powerful frameworks for sequential decision-making tasks. These models excel in scenarios requiring the progressive transformation of noisy inputs into structured outputs, enabling them to serve as effective controllers.
There are generally two diffusion-based frameworks for decision making. The first approach utilizes diffusion models to generate subsequent actions based on historical observations, effectively capturing temporal dependencies in sequential tasks [189]. On the other hand, the second one employs diffusion models to learn the trajectory of the decision-making process, modeling the progression of decisions over time and capturing the broader decision path.
Diffusion models can abstract the distribution of observation-action pairs as a policy. For example, diffusion Q-learning (Diffusion-QL) [156] employs a conditional diffusion model as its policy. This model learns an action-value function and seeks to optimize actions that align closely with the behavior policy by optimizing the loss function. The policy is trained by reducing the loss below:
@@
SfBC [158] decomposes the policy into two components: an expressive generative behavior model for capturing the data distribution and an action evaluation model for assessing action quality. DIPO [161] utilizes a diffusion policy to perform model-free online reinforcement learning, enabling efficient decision-making in complex environments.
Diffusion models also provide an effective way to describe the joint distribution of trajectories. Diffuser [153] treats trajectory optimization as a modeling problem, which is solved by iteratively refining trajectories using a diffusion probabilistic model. In this framework, a trajectory is represented as $\mathbf{\tau}=\begin{bmatrix}s_{0}&s_{1}&\cdots&s_{T}\\ a_{0}&a_{1}&\cdots&a_{T}\\ \end{bmatrix}$, where $T$ denotes the time horizon. During training, Diffuser minimizes a diffusion loss to optimize the trajectory generation process:
@@
Diffuser leverages classifier-guided sampling and image inpainting to reinterpret planning strategies. Decision Diffuser [154] models decision-making through conditional generative modeling, treating the policy as a return-conditioned diffusion model (see Fig. 10 for an overview). This approach allows Decision Diffuser to incorporate additional variables such as constraints and skills, enabling the generation of behaviors that meet multiple constraints simultaneously or demonstrate a combination of skills during test-time. AdaptDiffuser [160] creates rich synthetic expert data for goal-conditioned tasks by guiding the process with reward gradients and selecting high-quality data through a discriminator to finetune the diffusion model.
Diffusion models offer diverse applications in decision-making. Decision Stacks [155] deconstructs a goal-conditioned policy agent into three generative modules that use independent models trained in parallel with teacher forcing to simulate the temporal evolution of observations, rewards, and actions. Diffusion Policy [157] simulates robotic policy by converting a visuomotor policy into a denoising diffusion process conditioned on specific inputs. UniPi [159] casts sequential decision-making as a video generation task conditioned on text. The planner generates a sequence of future frames depicting planned actions based on a text-encoded goal, from which control actions are derived. Pearce et al. [152] attempts to imitate human behavior using diffusion models, abstracting a rich distribution over the joint action space.
@@ 4.1.6 GFlowNets as Controller
GFlowNet (GFN) is a powerful tool for sequential sampling, particularly in Markov decision processes (MDPs). In such a process, where a reward function $R(\cdot)$ is defined on each terminal state $s_{f}$, GFlowNet can be trained to learn a stochastic policy that determines the probability of any terminal state $s_{f}$ is proportional to $R(s_{f})$. The most basic form of GFlowNet is applied to discrete, acyclic MDP (DAGs), with both the state and action space are finite, and the states and their transitions form a directed acyclic graph (DAG). In this context, we define four key components of a GFlowNet:
  1. 1.
    An edge flow function $F(s,s^{\prime})$ which denotes the flow on the edge $s\rightarrow s^{\prime}$.
  2. 2.
    A vertex flow function $F(s)$:
    $F(s)=\sum_{s^{\prime}\in child(s)}F(s,s^{\prime}).$ (24)
  3. 3.
    Forward probability $P_{F}(s,s^{\prime})=\frac{F(s,s^{\prime})}{F(s)}$.
  4. 4.
    Backward probability $P_{B}(s,s^{\prime})=\frac{F(s^{\prime},s)}{F(s^{\prime})}$.
The induced policy of a GFlowNet is given by $P_{F}$,i.e., $\pi_{f(s,a)=s^{\prime}}(s,a)=P_{F}(s,s^{\prime})$.
GFlowNet [103, 51] is well-suited for complex scenarios where multiple trajectories may lead to the same terminal state. It transforms the set of trajectories into a flow and formulates the flow consistency equations as the learning objective, similar to how the Bellman equations are cast into Temporal Difference (TD) methods. GFlowNet demonstrates that any global minimum of the proposed objective results in a policy that samples from the desired distribution, achieving improved performance and diversity.
@@
Stochastic GFlowNet [168], an extension of GFlowNet, is tailored for stochastic environments. It decomposes state transitions into isolated environmental stochasticity, with a dynamics model trained to capture this stochasticity. GAFlowNets [169] introduces an intermediate reward as intrinsic motivation to address the exploration problem in sparse reward environments. This model combines edge-based and state-based intrinsic rewards to guide exploration effectively. Expected Flow Networks (EFlowNets) [170] adapts GFlowNets to stochastic environments as well, extending their applicability. Brunswic et al. [171] further generalizes GFlowNets to measurable spaces, including continuous state spaces, and extends the concept to handle cycles in this broader context.
CFlowNet [104] represents a significant advancement in the GFlowNet community by extending its capabilities to continuous control tasks (see Fig. 11). This extension is crucial as it bridges the gap between simplified simulation environments and complex real-world scenarios, expanding GFlowNet’s applicability to a broader range of practical applications. CFlowNet introduces a novel approach for solving continuous control tasks using GFlowNets. The authors present a theoretical formulation of CFlowNets and propose a unified framework that combines the action selection strategy, flow approximation method, and optimization objectives. Experimental results demonstrate that CFlowNets outperform many reinforcement learning methods, particularly in terms of exploration ability. A subsequent work, Continuous GFlowNet [106], develops a mathematical framework for continuous GFlowNet theory, further advancing the field.
@@ 4.1.7 Autoregressive Models as Controller
Autoregressive models excel in decision-making tasks by sequentially generating outputs conditioned on past observations. Decision Transformers (DTs) [176] advance this approach by defining reinforcement learning as conditional sequence modeling. It attempts to reformulate reinforcement learning as a conditional sequence modeling problem. It employs a causally masked Transformer to generate optimal actions aimed at maximizing the expected reward, conditioned on the past states, actions and the reward. The Online Decision Transformer [178] pretrains in the offline mode and fine-tunes online, unifying both within a framework that leverages sequence-level entropy regularizers and autoregressive modeling objectives to guide the learning process. DADT [190] improves the generalization ability by incorporating representation learning, which predicts the next state.
@@
Prompt-DT [191] extends DTs by introducing task-specific prompts as Fig. 13, enabling few-shot policy generalization and quick adaptation to new tasks, while maintaining the autoregressive decision-making framework.
Causal Decision Transformer for Recommender Systems (CDT4Rec) [192] applies DTs to recommendation systems by incorporating causal inference, predicting recommendations based on user history to maximize long-term reward. It serves as a controller, generating actions conditioned on past user interactions and system states.
Figure 13: Prompt-DT for few-shot policy generalization [191].
Trajectory Transformer [177] also extends this perspective by framing RL as sequence modeling over full trajectories. It models distributions over trajectories using a Transformer architecture and employs beam search for planning, enabling the generation of high-reward action sequences. BooT [182] further refines this approach by leveraging bootstrapping techniques and generating additional self-supervised offline data to enhance sequence model training.
Figure 14: The trajectory transformer trains on autoregressively discretized state, action, and reward sequences, using a planning process akin to language model sampling [177].
Pareto-Efficient Decision Agents (PEDA)[181] extends Decision Transformers [176] and RvS [193] for multi-objective reinforcement learning. It introduces a preference- and return-conditioned policy to handle multiple objectives efficiently in offline settings. With the advancement of large-scale language models [2], Transformer-based controllers have emerged as versatile solutions capable of handling a diverse set of tasks within a single unified model. GATO [179] introduces a generalist agent capable of handling multi-modal, multi-task, and multi-embodiment scenarios. This agent is versatile in executing a wide range of functions, including chatting, controlling robotic arms, and determining outputs such as text, joint torques, button presses, or other actions based on contextual input. Based on this work, RT-1 [194] develops a scalable model trained on a large-scale dataset of real-world tasks performed by real robots. It shows performance improvements that scale with data size, model capacity, and the diversity of tasks involved.

### v3

@@ 2.4 Difference with previous approaches
Reinforcement learning (RL) and generative models are distinct in their objectives and learning processes [102], as shown in Fig. 6. RL focuses on teaching an agent to learn an optimal policy through trial and error, maximizing cumulative rewards by interacting with the environment. The agent refines its strategy based on the feedback it receives from taking actions and observing the resulting states.
Generative models, on the other hand, aim to learn the underlying data distribution of a given dataset [103]. They generate new samples that resemble the training data, capturing the inherent patterns and structures.
The learning processes also differ. In RL, the agent explores the environment, takes actions, and receives rewards, continuously refining its policy to make more effective decisions. In contrast, generative models learn from a dataset and estimate the data’s probability distribution, with examples being VAEs and GANs.
Additionally, RL and generative models differ in their interaction with the environment [104]. In RL, the agent interacts with the environment through trial and error, learning from feedback without generating new data. In contrast, generative models, once trained, actively generate new data points that resemble the original data [105].
@@
In summary, RL focuses on learning optimal actions to achieve specific goals [106], while generative models aim to model data distributions and generate similar samples. Both are vital in machine learning but tackle different challenges.
@@
soft Q-learning [107] EBM Controller EBM Online RL Robot Control
@@
ALPINE [169] AM Modeler Decoder Only Online RL Optimization
ARP [170] AM Modeler Decoder Only Online RL Games
@@ 3 Taxonomy
Application. The last dimension highlights several representative applications where generative models have shown particular promise, including robot control, autonomous driving, gaming, structural generation, and optimization tasks.
@@ 4.1 Generative Models as Controller
A controller aims to find a policy that modifies the system’s inputs to attain the desired behavior [173, 174]. Generative models excel at creating control policies by modeling decision-making and generating diverse strategies. They learn from expert trajectories, capturing complex data distributions to extract optimal policies.
@@ 4.1.1 EBMs as Controller
EBMs serve as controllers by solving the Inverse Reinforcement Learning (IRL) problems, where the goal is to derive a control policy from expert demonstrations. In this context, the EBMs learn the underlying reward structure of the expert’s behavior. EBIL [108] refines the IRL process by estimating the expert’s energy with score matching, using it as a surrogate reward function to learn the policy through RL algorithms. Similarly, in [107], EBMs model policies in RL by incorporating an energy function that defines the desirability of actions in a given state. The goal is to minimize energy, guiding the agent toward optimal policies. Unlike traditional RL methods, this approach does not rely on predefined rewards but instead uses an energy-based framework for more flexible and scalable policy learning.
@@ 4.1.2 GANs as Controller
Generative Adversarial Imitation Learning (GAIL) [58] combines imitation learning with GAN to directly extract the policy from the data. It seeks to find a saddle point $(\pi,D)$, representing a pair consisting of a policy $\pi$ and a discriminator $D$, by optimizing the following expression:
@@
where $(s,a)$ denotes the state-action pair, $\pi$ is the policy, $\pi_{E}$ represents the expert policy and $H$ is a regularizer.
Several extensions have been proposed to address GAIL’s limitations. InfoGAIL [112] extends GAIL by unsupervisedly learning the structure of expert demonstrations, allowing it to imitate complex behaviors and extract interpretable representations. FAGIL [114], another extension, adds a safety layer to ensure policy safety while enabling closed-form computation of densities and gradients, combining adversarial training with worst-case safety guarantees for both safety and performance.
@@
Other related approaches include Adversarial Inverse Reinforcement Learning (AIRL) [117], which frames policy learning as an adversarial reward learning problem for diverse environments. AugAIRL [118] enhances AIRL by incorporating semantic rewards, while WAIL [119] connects inverse reinforcement learning with optimal transport for large-scale applications. Model-based AIRL (MAIRL) [120] improves policy optimization by using a self-attention dynamics model for differentiable computation, and SC-AIRL [175] reduces exploration by breaking long-horizon tasks into subtasks with shared rewards and critics.
@@ 4.1.3 VAEs as Controller
VAEs serve as controllers by learning compact, structured latent spaces that encode high-dimensional environments and decision-making processes, which are directly used to generate control actions.
For example, Play-LMP [128] uses a self-supervised method to learn control from play behaviors, allowing autonomous policy generation. TACO-RL [129] extends this by adopting a hierarchical model for long-horizon policies from high-dimensional camera observations. OPAL [130] applies VAE for offline learning of primitive actions in imitation tasks. GTI [126] uses VAE in a two-stage imitation learning algorithm to improve generalization in complex environments. HULC [127] enhances performance by integrating hierarchical decomposition in robot control with a multimodal transformer encoder and discrete latent plans.
MaskDP [131] introduces masked autoencoding to enhance decision-making models by learning latent spaces in a scalable, generalizable way, improving performance in high-dimensional tasks. This approach, similar to VAEs, generates control policies through efficient latent representations while incorporating masking for better task generalization.
@@ 4.1.4 Normalizing Flows as Controller
Normalizing flows are gaining attention for their ability to learn complex probability distributions and perform efficient density estimation. While mainly used in generative modeling, they are also appliable to decision-making control tasks [67, 81]. In this context, normalizing flows can model and approximate the decision space underlying [47, 79].
A common application of normalizing flows for controllers is behavioral cloning, where the target distribution $\mu$ represents the behavior to replicate. By training a reversed normalizing flow, where $\nu$ is a complex distribution with defined density and $\mu$ is a simpler distribution like a normal distribution, the training process becomes feasible.
Normalizing Flow Policy (NF Policy) [135] integrates Normalizing Flows with the Soft Actor-Critic (SAC) framework, enabling more expressive policy learning. Conservative Normalizing Flows (CNF) [136] uses NFs as a conservative action encoder for learning policies in latent action spaces. The action encoder is first pre-trained on an offline dataset, followed by policy learning through reinforcement learning. Guided Flows [137] generates plans in offline reinforcement learning by training Continuous Normalizing Flows with classifier-free guidance to regress vector fields.
@@ 4.1.5 Diffusion Models as Controller
Diffusion models, known for their iterative refinement and controlled noise perturbation processes, have shown to be effective in sequential decision-making tasks. These models excel in transforming noisy inputs into structured outputs, making them well-suited as controllers.
There are two main diffusion-based frameworks for decision-making. The first generates subsequent actions based on historical observations, capturing temporal dependencies in sequential tasks [176]. The second models the entire trajectory of the decision-making process, capturing the broader decision path over time.
Diffusion models can represent the distribution of observation-action pairs as a policy. For example, Diffusion Q-learning (Diffusion-QL) [143] uses a conditional diffusion model as its policy. It learns an action-value function and optimizes actions by minimizing the following loss function:
@@
SfBC [145] breaks the policy into two components: a generative behavior model to capture the data distribution and an action evaluation model to assess action quality. DIPO [148] applies a diffusion policy for model-free, online reinforcement learning in complex environments.
Diffusion models can also describe the joint distribution of trajectories. Diffuser [140] optimizes trajectories through iterative refinement with a diffusion probabilistic model. In this framework, a trajectory is represented as $\mathbf{\tau}=\begin{bmatrix}s_{0}&s_{1}&\cdots&s_{T}\\ a_{0}&a_{1}&\cdots&a_{T}\\ \end{bmatrix}$, where $T$ denotes the time horizon. During training, Diffuser minimizes a diffusion loss to optimize the trajectory generation process:
@@
Diffuser uses classifier-guided sampling and image inpainting for planning strategies. Decision Diffuser [141] treats decision-making as conditional generative modeling, using a return-conditioned diffusion model (see Fig. 8). This approach allows for incorporating constraints and skills into the generated behaviors. AdaptDiffuser [147] uses reward gradients and a discriminator to guide synthetic expert data generation for goal-conditioned tasks, further finetuning the diffusion model.
Diffusion models are applied in diverse decision-making contexts. Decision Stacks [142] deconstructs goal-conditioned policies into three generative modules that simulate the evolution of observations, rewards, and actions in parallel. Diffusion Policy [144] simulates robotic policies by converting visuomotor policies into a denoising diffusion process conditioned on inputs. UniPi [146] frames sequential decision-making as a video generation task conditioned on text, where control actions are derived from a text-encoded goal. Pearce et al. [139] uses diffusion models to imitate human behavior, abstracting the distribution over the joint action space.
@@ 4.1.6 GFlowNets as Controller
GFlowNet (GFN) is a powerful tool for sequential sampling, particularly in Markov Decision Processes (MDPs). It learns a stochastic policy where the probability of a terminal state $s_{f}$ is proportional to a reward function $R(s_{f})$. GFlowNet is typically applied to discrete MDPs, but it is not limited to acyclic structures and can also be used in other types of MDPs, including those with more general graph structures.
GFlowNet [91, 45] is suitable for scenarios with multiple trajectories leading to the same terminal state. It transforms these trajectories into a flow and uses flow consistency equations as the learning objective, similar to the Bellman equations in Temporal Difference (TD) methods. A global minimum of this objective results in a policy that samples from the desired distribution, improving performance and diversity. Stochastic GFlowNet [155] extends GFlowNet to stochastic environments, capturing environmental randomness. GAFlowNets [156] adds intermediate rewards to address exploration challenges in sparse reward environments. EFlowNets [157] further adapts GFlowNets for such environments. Brunswic et al. [93] generalize GFlowNets to measurable spaces, including continuous state spaces and cycles.
@@
CFlowNet [92] advances GFlowNet by extending its use to continuous control tasks (see Fig. 9), making it applicable to real-world scenarios. It introduces a unified framework combining action selection, flow approximation, and optimization objectives. Experimental results show CFlowNets outperform traditional reinforcement learning methods, especially in exploration. Continuous GFlowNet [94] further develops the theoretical framework for continuous control.
@@ 4.1.7 Autoregressive Models as Controller
Autoregressive models are effective in decision-making tasks by generating outputs sequentially based on past observations. Decision Transformers (DTs) [162] reformulate RL as a conditional sequence modeling problem. They use a causally masked Transformer to generate optimal actions that maximize expected rewards, conditioned on past states, actions, and rewards. The Online Decision Transformer [164] combines offline pretraining with online fine-tuning, using sequence-level entropy regularizers and autoregressive objectives. DADT [177] enhances generalization by predicting the next state through representation learning. Trajectory Transformer [163] extends this idea by modeling reinforcement learning as sequence modeling over entire trajectories. It uses a Transformer to model trajectory distributions and employs beam search for planning to generate high-reward sequences. BooT [168] further enhances this by generating self-supervised offline data to improve model training.
Pareto-Efficient Decision Agents (PEDA) [167] extends Decision Transformers [162] and RvS [178] for multi-objective reinforcement learning, introducing a preference- and return-conditioned policy for efficient multi-objective handling. With the rise of large-scale language models [2], Transformer-based controllers, such as GATO [165], have emerged as versatile agents capable of handling multi-modal, multi-task, and multi-embodiment scenarios. RT-1 [179] builds on this by training a scalable model on real-world tasks, demonstrating performance improvements with larger data and model sizes.